# PlantTraitNet: An Uncertainty-Aware Multimodal Framework for Global-Scale Plant Trait Inference from Citizen Science Data

### v1

@@ Results
Table 1: Global trait map benchmarking against sPlotOpen CWMs (1° resolution). Best. Second-best. External products: Schiller (schiller2021deep), Wolf (wolf2022citizen), Moreno (moreno2018methodology), Butler (butler2017mapping), Boonman (boonman2020assessing), Madani (madani2018future), Van Bodegom (van2014fully).
@@
Multi-Task
DinoV2 0.15 ± 0.00 0.31 ± 0.00 0.32 ± 0.00 0.14 ± 0.01 1
@@
We benchmark our global trait maps derived from models trained on both raw and filtered data and those of previous studies using R², Pearson’s $r$, and nMAE (Table 1). Depending on the metric, our model for LN delivers comparable results with those by Schiller et al. (schiller2021deep). For the other three traits, our models consistently achieve higher performance than previously published plant trait. The overall improvement reflects the model’s ability to capture complex and variable patterns in large-scale trait prediction. While the $r$ scores suggest that the maps capture relative differences, substantially lower R² scores indicate that PlantTraitNet maps and all other products are systematically biased (also see Appendix), underscoring the inherent challenge of revealing morphological and physiological ecosystem patterns at global scale.
@@
To reveal trait variability within the biosphere on a global scale, a computer vision model must be robust across species and thus phylogenetic lineages. Using the inferences and species information of the validation data, we show that residuals from PlantTraitNet are largely unsystematically distributed in phylogenetic space (Fig. 5). To quantify this, we use two standard metrics of phylogenetic signal:
Pagel’s $\lambda$, which captures broad-scale phylogenetic autocorrelation in residual covariance (pagel1999inferring) and Blomberg’s $K$ is, which is more sensitive to fine-scale signal among closely related species (blomberg2003testing). For SLA and leaf nitrogen, the phylogenetic signal is weak ($\lambda$ = 0.04 and 0.15; $K$ = 0.0053 and 0.0076), suggesting that prediction errors are largely independent of species relatedness. Although errors for height ($\lambda$ = 0.80, $K$ = 0.018) and leaf area ($\lambda$ = 0.56, $K$ = 0.0067) show some phylogenetic autocorrelation, the consistently low $K$ values indicate that even closely related species do not share systematic prediction biases (see Appendix for details). Although PlantTraitNet was trained using weak annotations on a species level, these findings underscore the model’s strong generalizability and robustness across the plant tree of life.
Despite weak supervision at the species level, PlantTraitNet captures within-species variability in trait expression. This is particularly evident in the case of height prediction (Fig. 6), where the model reflects differences across growth forms and developmental stages within individual species (see Appendix for within-species variability for the traits). This suggests that the model is not simply regressing to a species-level mean but is sensitive to morphological cues in the images that reflect ecological and ontogenetic variation.
@@
To evaluate the contribution of each input modality to trait prediction, we conduct an ablation study using different combinations of image, geospatial, and depth information (Table 2). Our goal is to understand how each modality influences model performance across key plant functional traits: H, LA, SLA, and LN. For this ablation study, we also experimented with a pretrained BioCLIP (stevens2024bioclip) encoder as an alternative to DinoV2. For BioCLIP, we extracted the embedding from its classification token, as it empirically showed superior performance (detailed in the Appendix). For geospatial priors, we assess SatCLIP (klemmer2025satclip) (trained on satellite imagery) for its ability to capture vegetation density and phenology, GeoCLIP (vivanco2023geoclip) (a geo-localization model trained on natural images), and Climplicit (dollinger2025climplicit).
@@ Discussion
Predicting plant traits from citizen science photos is challenging due to data variability and biases, including spatial and taxonomic bias, and overrepresentation of smaller growth forms like grasses and herbs (di2021observing; sierra2024divshift). Ecological complexity adds difficulty, as traits vary across biomes, with generalists showing common traits and specialists distinct ones, resulting in skewed, long-tailed distributions (see Appendix). Unlike animals with fixed body plans, e.g. with symmetric and fixed numbers of legs or arms, plants have a comparably flexible morphology, resulting in varying numbers of plant organs, such as leaves or branches. This variability complicates trait prediction via computer vision. Despite these challenges, our results show promising potential. Future work should focus on reducing biases through targeted data acquisition. Increased acquisition of reference data to enable better ‘turning point’ selection and incorporate label correction to enhance model robustness and generalization in ecological contexts.
@@ Acknowledgements
The preparation of our citizen science training data involved several preprocessing steps to ensure quality and ecological representativeness. Using the geolocation metadata of the images, we removed samples from aquatic and urban areas based on the ESA WorldCover 10m v100 land cover map (zanaga2021esa).
We weakly annotate these images with species-level trait values from the TRY database (kattge2020try). To reduce label noise due to sparsely represented species, we excluded species with fewer than three trait observations in the TRY database. Additionally, to further mitigate the impact of outliers, we computed the species-level median trait values after removing observations below the 5th and above the 99th percentiles.
Finally, we constructed the training and validation splits using an 80-20 partition, stratified by plant growth form (trees, shrubs, grasses) to preserve diversity across the subsets. Information on plant growth form was derived from the TRY database (kattge2020try, Trait-ID 4). For each species, we standardized the provided growth forms into 3 classes (Tree, Shrub, Grassland) and subsequently applied a majority vote per species (lusk2025smartphones).
@@
Our uncertainty-guided data cleaning strategy is grounded in the concept of a turning point—a stage during training when the model transitions from learning generalizable patterns to overfitting noisy labels (lu2022selc). Detecting this turning point is critical for effective uncertainty-aware data cleaning. However, in weakly supervised settings, where ground truth labels are absent or imprecise, identifying this transition reliably is challenging. To overcome this limitation, we curated a high-quality reference dataset consisting of plant images paired with accurate trait measurements. This dataset consists of smaller datasets distributed globally from various previous studies and a range of collaborators. Each collaborator contributed plant trait measurements and associated images from ecologically diverse locations, including Germany, La Palma, India, Australia, Panama, Canada, Indonesia, Switzerland, Portugal, and Namibia (see Figure 8). Detailed statistics on sample sizes and taxonomic diversity, quantified as the number of unique species, are reported in Table 3. Because different collaborators focused on specific traits, we maintain separate datasets per trait.
@@ Results
Figure 12 show global maps at 1 degree resolution obtained from aggregating predicted traits from PlantTraitNet on 300k citizen science images. In addition to the quantitative evaluation against vegetation surveys (sPlotOpen, next section), we found that the global patterns reproduce expected global trait patterns (lusk2025smartphones; wolf2022citizen). For instance, for height and leaf area, we find expected longitudinal variation, such as large plants with big leaves towards the equator. Most robust plants (high specific leaf area and nitrogen content) are found in temperate and boreal zones.
@@
This analysis provides insight into how well the model reflects trait variability within species, relative to aggregated trait observations. Our results, visualized in Figures 15, 16, 17 and 18, consistently show that the model learns to represent a wide range of intraspecific variation for traits like Height (H) and Leaf Area (LA). In these cases, the predicted trait ranges often exceeded the variation captured in the original TRY database observations. This seems plausible, given that trait observations are typically performed on healthy and adult plants. Compared to Leaf Area (LA) and Plant Height (H), for physiological traits such as Specific Leaf Area (SLA) and Leaf Nitrogen (LN), we observe little predicted variation. We hypothesize this is a positive outcome, as these leaf traits are ecologically expected to show less intraspecific variability on the premise that interspecific trait variation generally exceeds intraspecific variation (dong2020components; wright2017global). For improved visibility, only predicted values between the 5th and 95th percentiles are shown.
@@
A phylogenetic tree was constructed from hierarchical taxonomic information (Kingdom to Species) using as.phylo() from the ape package (paradis2019ape). Mean relative errors (MRE) were computed per species, normalized by the trait-specific error range, and matched with taxonomic metadata. In Figure 14, MREs are shown at the family and class levels, with no clear visual pattern of systematic bias. Moreover, we report standard metrics to quantify the phylogenetic signal in species-level prediction errors, calculated via the phyloSignal() function from the phylosignal R package. Pagel’s $\boldsymbol{\lambda}$ (pagel1999inferring) measures the overall fit of the data to the phylogeny by scaling internal branches; values close to 1 indicate a strong phylogenetic structure, while values near 0 suggest independence from phylogeny. Statistical significance is assessed by comparing the fitted $\lambda$ to a model with $\lambda=0$. Blomberg’s $\boldsymbol{K}$ (blomberg2003testing) compares the observed trait variance across the tree to expectations under a Brownian motion model; values close to 1 indicate strong signal, and values $\ll$1 suggest weak signal, especially among closely related species. $\boldsymbol{K^{*}}$ is a variation of $K$ that is less sensitive to tree imbalance and branch length distortions. Together, these metrics offer complementary perspectives on phylogenetic structure at different evolutionary scales (see Table 5). Pagel’s $\lambda$ captures broad-scale phylogenetic autocorrelation in residual covariance (pagel1999inferring), while Blomberg’s $K$ is more sensitive to fine-scale signal among closely related species (blomberg2003testing). For SLA and leaf nitrogen, the phylogenetic signal is weak ($\lambda$ = 0.04 and 0.15; $K$ = 0.0053 and 0.0076), suggesting that prediction errors are largely independent of species relatedness. Although errors for height ($\lambda$ = 0.80, $K$ = 0.018) and leaf area ($\lambda$ = 0.56, $K$ = 0.0067) show some phylogenetic autocorrelation, the consistently low $K$ values indicate that even closely related species do not share systematic prediction biases. These results reveal that while slight patterns may emerge at the level of broader plant classes, there is no consistent phylogenetic structure among closely related species. Altogether, this underscores that despite being trained on species-level annotations, the model exhibits strong generalizability and methodological robustness for predicting plant traits across the plant kingdom.
@@
While the classification token (‘[CLS]’) is a common choice for a global image representation, the individual patch tokens of vision transformers contain rich, spatially-aware information that can be more beneficial for downstream tasks (caron2021emerging). We therefore perform an ablation study to determine the optimal token representation for our trait prediction task (Table 7).
@@
We observe that using a 768-dimensional image embedding from DINOv2, a 256-dimensional geolocation embedding, and a 768-dimensional multimodal fusion backbone tends to yield strong performance across traits. This trend is consistent across the tested geolocation encoders: Climplicit (dollinger2025climplicit), SatCLIP (klemmer2025satclip), and GeoCLIP (vivanco2023geoclip). Among these, Climplicit performs best in our experimental setup.

### v2

@@ Results
Table 1: Global trait map benchmarking against sPlotOpen CWMs (1° resolution). Best. Second-best. External products: Schiller (Schiller et al. 2021), Wolf (Wolf et al. 2022), Moreno (Moreno-Martínez et al. 2018), Butler (Butler et al. 2017), Boonman (Boonman et al. 2020), Madani (Madani et al. 2018), Bodegom (Van Bodegom et al. 2014). PlantTraitNet (Ours) is evaluated after training on both raw and refined datasets.
@@
Multi-Task DinoV2 0.15 ± 0.00 0.31 ± 0.00 0.32 ± 0.00 0.14 ± 0.01 1
@@
We benchmark our global trait maps derived from models trained on both raw and filtered data and those of previous studies using R², Pearson’s $r$, and nMAE (Table 1). Depending on the metric, our model for LN delivers comparable results with those by Schiller et al. (Schiller et al. 2021). For the other three traits, our models consistently achieve higher performance than previously published plant trait maps. The overall improvement reflects the model’s ability to capture complex and variable patterns in large-scale trait prediction. While the $r$ scores suggest that the maps capture relative differences, substantially lower R² scores indicate that PlantTraitNet maps and all other products are systematically biased (also see Appendix), underscoring the inherent challenge of revealing morphological and physiological ecosystem patterns at global scale.
@@
To reveal trait variability within the biosphere on a global scale, a computer vision model must be robust across species, and thus phylogenetic lineages. Using inferences and species information from the validation data, we show that the residuals of PlantTraitNet are largely unsystematically distributed in the phylogenetic space (Fig. 5). To quantify this, we use two standard metrics of phylogenetic signal:
Pagel’s $\lambda$, which captures broad-scale phylogenetic autocorrelation in residual covariance (Pagel 1999) and Blomberg’s $K$, which is more sensitive to fine-scale signals among closely related species (Blomberg et al. 2003). For SLA and leaf nitrogen, the phylogenetic signal is weak ($\lambda$ = 0.04 and 0.15; $K$ = 0.0053 and 0.0076), suggesting that prediction errors are largely independent of species relatedness. Although errors for height ($\lambda$ = 0.80, $K$ = 0.018) and leaf area ($\lambda$ = 0.56, $K$ = 0.0067) show some phylogenetic autocorrelation, the consistently low $K$ values indicate that even closely related species do not share systematic prediction biases (see Appendix for details). Although PlantTraitNet was trained using weak annotations at the species level, these findings underscore the model’s strong generalizability and robustness across the plant tree of life.
Despite weak supervision at the species level, PlantTraitNet captures within-species variability in trait expression. This is particularly evident in the case of height prediction (Fig. 6), where the model reflects differences across growth forms and developmental stages within individual species (see Appendix for additional examples of within-species variability across traits). This suggests that the model is not simply regressing to a species-level mean but is sensitive to morphological cues in the images that reflect ecological and ontogenetic variation.
@@
To evaluate the contribution of each input modality to trait prediction, we conduct an ablation study using different combinations of image, geospatial, and depth information (Table 2). Our goal is to understand how each modality influences model performance across key plant functional traits: H, LA, SLA, and LN. For this ablation study, we also experimented with a pretrained BioCLIP (Stevens et al. 2024) encoder as an alternative to DinoV2. For BioCLIP, we extracted the embedding from its classification token, as it empirically showed superior performance (detailed in the Appendix). For geospatial priors, we assess SatCLIP (Klemmer et al. 2025), which is trained on satellite imagery and captures vegetation density and phenology; GeoCLIP (Vivanco Cepeda et al. 2023), a geo-localization model trained on natural images; and Climplicit  (Dollinger et al. 2025)
@@ Discussion
Predicting plant traits from citizen science photos is challenging due to data variability and biases, including spatial and taxonomic bias, and overrepresentation of smaller growth forms such as grasses and herbs (Di Cecco et al. 2021; Sierra et al. 2024). Ecological complexity adds difficulty, as traits vary across biomes, with generalists showing common traits and specialists distinct ones, resulting in skewed, long-tailed distributions (see Appendix). Unlike animals with fixed body plans, e.g. with symmetric and fixed numbers of legs or arms, plants have a comparably flexible morphology, resulting in varying numbers of plant organs, such as leaves or branches. This variability complicates trait prediction via computer vision. Despite these challenges, our results show promising potential. Future work should focus on reducing biases through targeted data acquisition. Increased acquisition of reference data to enable better ‘turning point’ selection and incorporate label correction to improve model robustness and generalization in ecological contexts.
@@ Acknowledgements
The preparation of our citizen science training data involved several preprocessing steps to ensure quality and ecological representativeness. Using the geolocation metadata of the images, we removed samples from aquatic and urban areas based on the ESA WorldCover 10m v100 land cover map (Zanaga et al. 2021).
We weakly annotate these images with species-level trait values from the TRY database (Kattge et al. 2020). To reduce label noise due to sparsely represented species, we excluded species with fewer than three trait observations in the TRY database. Additionally, to further mitigate the impact of outliers, we computed the species-level median trait values after removing observations below the 5th and above the 99th percentiles.
Finally, we constructed the training and validation splits using an 80-20 partition, stratified by plant growth form (trees, shrubs, grasses) to preserve diversity across the subsets. Information on plant growth form was derived from the TRY database (Kattge et al. 2020, Trait-ID 4). For each species, we standardized the provided growth forms into 3 classes (Tree, Shrub, Grassland) and subsequently applied a majority vote per species (Lusk et al. 2025).
@@
Our uncertainty-guided data cleaning strategy is grounded in the concept of a turning point—a stage during training when the model transitions from learning generalizable patterns to overfitting noisy labels (Lu and He 2022). Detecting this turning point is critical for effective uncertainty-aware data cleaning. However, in weakly supervised settings, where ground truth labels are absent or imprecise, identifying this transition reliably is challenging. To overcome this limitation, we curated a high-quality reference dataset consisting of plant images paired with accurate trait measurements. This dataset consists of smaller datasets distributed globally from various previous studies and a range of collaborators. Each collaborator contributed plant trait measurements and associated images from ecologically diverse locations, including Germany, La Palma, India, Australia, Panama, Canada, Indonesia, Switzerland, Portugal, and Namibia (see Figure 8). Detailed statistics on sample sizes and taxonomic diversity, quantified as the number of unique species, are reported in Table 3. Because different collaborators focused on specific traits, we maintain separate datasets per trait.
@@ Results
Figure 12 show global maps at 1 degree resolution obtained from aggregating predicted traits from PlantTraitNet on 300k citizen science images. In addition to the quantitative evaluation against vegetation surveys (sPlotOpen, next section), we found that the global patterns reproduce expected global trait patterns (Lusk et al. 2025; Wolf et al. 2022). For instance, for height and leaf area, we find expected longitudinal variation, such as large plants with big leaves towards the equator. Most robust plants (high specific leaf area and nitrogen content) are found in temperate and boreal zones.
@@
This analysis provides insight into how well the model reflects trait variability within species, relative to aggregated trait observations. Our results, visualized in Figures 15, 16, 17 and 18, consistently show that the model learns to represent a wide range of intraspecific variation for traits like Height (H) and Leaf Area (LA). In these cases, the predicted trait ranges often exceeded the variation captured in the original TRY database observations. This seems plausible, given that trait observations are typically performed on healthy and adult plants. Compared to Leaf Area (LA) and Plant Height (H), for physiological traits such as Specific Leaf Area (SLA) and Leaf Nitrogen (LN), we observe little predicted variation. We hypothesize this is a positive outcome, as these leaf traits are ecologically expected to show less intraspecific variability on the premise that interspecific trait variation generally exceeds intraspecific variation (Dong et al. 2020; Wright et al. 2017). For improved visibility, only predicted values between the 5th and 95th percentiles are shown.
@@
A phylogenetic tree was constructed from hierarchical taxonomic information (Kingdom to Species) using as.phylo() from the ape package (Paradis and Schliep 2019). Mean relative errors (MRE) were computed per species, normalized by the trait-specific error range, and matched with taxonomic metadata. In Figure 14, MREs are shown at the family and class levels, with no clear visual pattern of systematic bias. Moreover, we report standard metrics to quantify the phylogenetic signal in species-level prediction errors, calculated via the phyloSignal() function from the phylosignal R package. Pagel’s $\boldsymbol{\lambda}$ (Pagel 1999) measures the overall fit of the data to the phylogeny by scaling internal branches; values close to 1 indicate a strong phylogenetic structure, while values near 0 suggest independence from phylogeny. Statistical significance is assessed by comparing the fitted $\lambda$ to a model with $\lambda=0$. Blomberg’s $\boldsymbol{K}$ (Blomberg et al. 2003) compares the observed trait variance across the tree to expectations under a Brownian motion model; values close to 1 indicate strong signal, and values $\ll$1 suggest weak signal, especially among closely related species. $\boldsymbol{K^{*}}$ is a variation of $K$ that is less sensitive to tree imbalance and branch length distortions. Together, these metrics offer complementary perspectives on phylogenetic structure at different evolutionary scales (see Table 5). Pagel’s $\lambda$ captures broad-scale phylogenetic autocorrelation in residual covariance (Pagel 1999), while Blomberg’s $K$ is more sensitive to fine-scale signal among closely related species (Blomberg et al. 2003). For SLA and leaf nitrogen, the phylogenetic signal is weak ($\lambda$ = 0.04 and 0.15; $K$ = 0.0053 and 0.0076), suggesting that prediction errors are largely independent of species relatedness. Although errors for height ($\lambda$ = 0.80, $K$ = 0.018) and leaf area ($\lambda$ = 0.56, $K$ = 0.0067) show some phylogenetic autocorrelation, the consistently low $K$ values indicate that even closely related species do not share systematic prediction biases. These results reveal that while slight patterns may emerge at the level of broader plant classes, there is no consistent phylogenetic structure among closely related species. Altogether, this underscores that despite being trained on species-level annotations, the model exhibits strong generalizability and methodological robustness for predicting plant traits across the plant kingdom.
@@
While the classification token (‘[CLS]’) is a common choice for a global image representation, the individual patch tokens of vision transformers contain rich, spatially-aware information that can be more beneficial for downstream tasks (Caron et al. 2021). We therefore perform an ablation study to determine the optimal token representation for our trait prediction task (Table 7).
@@
We observe that using a 768-dimensional image embedding from DINOv2, a 256-dimensional geolocation embedding, and a 768-dimensional multimodal fusion backbone tends to yield strong performance across traits. This trend is consistent across the tested geolocation encoders: Climplicit (Dollinger et al. 2025), SatCLIP (Klemmer et al. 2025), and GeoCLIP (Vivanco Cepeda et al. 2023). Among these, Climplicit performs best in our experimental setup.